# Topology and Spectrum in Measurement-Induced Phase Transitions

### v1

@@ Abstract
We consider a $(1+1)$-dimensional quantum circuit acting on $2L$ Majorana fermions [see Fig. 1(a)], which consists of repeated applications of three elementary steps: (i) unitary time evolution $\hat{\mathcal{U}}$, (ii) measurements of all Majorana pairs on odd bonds, and (iii) measurements of all Majorana pairs on even bonds. Applying these three operations on a state |$\ket{\psi_{t}}$, we obtain a state |$\ket{\psi_{t+1}}$ updated by a single time step. To be more precise, the unitary evolution is given by the Kitaev chain Hamiltonian, $\hat{\mathcal{U}}=e^{-i\hat{\mathcal{H}}_{\mathrm{Kitaev}}}$, with
@@
which are weak measurements of the strength $\theta_{e/o}=\tanh^{-1}(\mu_{e/o})$. We set $\mu_{e}=1-\mu_{o}$ hereafter. Given a (unnormalized) state |$\ket{\psi}$, each outcome $s$ is obtained with the Born probability, $p^{e/o}_{j}(s)=\bra{\psi}(\hat{\mathcal{K}}_{j}^{e/o}(s))^{\dagger}\hat{\mathcal{K}}_{j}^{e/o}(s)\ket{\psi}/{\langle\psi|\psi\rangle}.$ In the following, we use three boundary conditions: open (OBC), periodic (PBC), and antiperiodic (APBC). For the unitary evolution $\hat{\mathcal{U}}$, the boundary condition is simply specified by $J^{\prime}$: $J^{\prime}=0,J,$ and $-J$ for the OBC, PBC, and APBC, respectively. For the measurements, the OBC means that we discard $\hat{\mathcal{K}}^{e}_{L}(s)$ from the circuit, while we need a careful definition of the APBC, as we will explain later.
@@
Since both unitary evolution and measurements are bilinear in Majorana fermions, this circuit maps a fermionic Gaussian state to another Gaussian state [46, 47, 48, 49, 50, 51, 52]. A fermionic Gaussian state is completely characterized by the Majorana covariance matrix $\Gamma_{t}$ whose element reads $(\Gamma_{t})_{\ell\ell^{\prime}}=(i/2)\bra{\psi_{t}}[\gamma_{\ell},\gamma_{\ell^{\prime}}]\ket{\psi_{t}}/\langle\psi_{t}|\psi_{t}\rangle$. Time evolution of the (unnormalized) Gaussian state is encoded in the transformation of the Majorana fermions, $\vec{\gamma}\to\hat{\mathcal{K}}_{T}(\bm{s})^{\dagger}\vec{\gamma}(\hat{\mathcal{K}}_{T}(\bm{s})^{\dagger})^{-1}=K_{T}(\bm{s})\vec{\gamma}$, where
$K_{T}(\bm{s})=\prod_{t=1}^{T}\left(e^{-i\tilde{\Theta}^{e}(\bm{s}_{t})}e^{-i\tilde{\Theta}^{o}(\bm{s}_{t})}e^{\tilde{H}_{\mathrm{Kitaev}}}\right).$ (4)
Here, $\tilde{H}_{\mathrm{Kitaev}}$ and $\tilde{\Theta}^{e/o}(\bm{s}_{t})$ are real antisymmetric matrices defined through ^$\hat{\mathcal{H}}_{\mathrm{Kitaev}}{={i}\vec{\gamma}^{\mathsf{T}}\tilde{H}_{\mathrm{Kitaev}}\vec{\gamma}}/4$ and $\prod_{j}\hat{\mathcal{K}}^{e/o}_{j}(s_{2j/2j-1,t})\propto\exp({i\vec{\gamma}^{\mathsf{T}}\tilde{\Theta}^{e/o}(\bm{s}_{t})\vec{\gamma}/4})$ with $\vec{\gamma}=(\gamma_{1},\cdots,\gamma_{2L})^{\mathsf{T}}$ (see Supplemental Material [53]).
@@
$H_{\textrm{eff},T}(\bm{s})=-\frac{i}{2T}\ln[K_{T}(\bm{s})K_{T}(\bm{s})^{\dagger}],$ (5)
and the corresponding many-body Hamiltonian ^$\hat{\mathcal{H}}_{\textrm{eff},T}(\bm{s})=-i\vec{\gamma}^{\mathsf{T}}H_{\textrm{eff},T}(\bm{s})\vec{\gamma}/4$. Since $H_{\textrm{eff},T}(\bm{s})$ is real antisymmetric, its eigenvalues $iz_{\ell,T}(\bm{s})\in i\mathbb{R}$ come in pairs, $z_{2j-1,T}(\bm{s})=-z_{2j,T}(\bm{s})$. This defines the single-particle energy spectrum $z_{\ell,T}(\bm{s})$, which in the asymptotic limit coincides with the Lyapunov spectrum: $z_{\ell}=\lim_{T\to\infty}z_{\ell,T}(\bm{s})$. We below arrange them as $0\leq z_{1}=-z_{2}\leq\cdots\leq z_{2L-1}=-z_{2L}$. We can also compute the corresponding Lyapunov vectors $\vec{w}_{\ell,T}(\bm{s})$. Written in the matrix form $\tilde{O}_{T}(\bm{s})=(\vec{w}_{1,T}(\bm{s}),\cdots,\vec{w}_{2L,T}(\bm{s}))$, they give in the asymptotic limit an orthogonal matrix that brings $H_{\textrm{eff},T}(\bm{s})$ into the standard form,
$\lim_{T\to\infty}\tilde{O}_{T}(\bm{s})^{\mathsf{T}}H_{\textrm{eff},T}(\bm{s})\tilde{O}_{T}(\bm{s})=\bigoplus_{j=1}^{L}\begin{pmatrix}{0}&z_{2j-1}\\ -z_{2j-1}&{0}\end{pmatrix}.$ (6)
In practice, we construct $\tilde{O}_{T}(\bm{s})$ through computing Lyapunov vectors corresponding to the non-negative Lyapunov spectrum based on the complex fermion representation, which is equivalent to the Majorana representation [53]. In the following, we only look at a single trajectory specified by a typical sequence of $\bm{s}$ and consider a temporal average of quantities in the long-time limit unless otherwise mentioned [53].
@@
Figure 2: (a) Non-negative single-particle Lyapunov spectra for $L=64$ and (b) sum of squared amplitudes of the spinor wavefunction corresponding to the lowest energy $z_{1}$ at the edges for the measurement-only circuit with the OBC. We find that the topologically non-trivial phase hosts a spatially localized Majorana edge mode with the zero Lyapunov exponent. The inset in (b) shows the spatial distribution of squared amplitudes of the spinor wavefunction at $t=2L$ averaged over $1000$ different trajectories for $L=16$ and $\mu_{e}=0.1$.
@@
Remarkably, we reveal that the topologically nontrivial area-law phase is characterized by Majorana zero modes for the Lyapunov spectrum. In Fig. 2(a), we plot the non-negative single-particle Lyapunov spectrum $z_{k}$ against $\mu_{e}$ for the measurement-only circuit under the OBC for $L=64$. Importantly, $z_{1}$ takes a finite value for $\mu_{e}>0.5$, while it vanishes for $\mu_{e}\leq 0.5$. Vanishing of $z_{1}$ under the OBC in the topological area-law phase implies the many-body Lyapunov gap closing for ^$\hat{\mathcal{H}}_{\textrm{eff},T}(\bm{s})$, which is reminiscent of Majorana zero modes for the static Kitaev chain [64].
@@
$P^{\mathrm{PBC/APBC}}_{T}(\bm{s})$ $=\det[\tilde{O}^{\textrm{PBC/APBC}}_{T}(\bm{s})].$ (9)
Since $\chi_{T}(\bm{s})$ approximates and converges to $Q_{T}(\bm{s})$ for $T\to\infty$, we will also call $\chi_{T}(\bm{s})$ as the topological invariant. We note that $\tilde{O}^{\textrm{PBC/APBC}}_{T}(\bm{s})$ is not an orthogonal matrix for general $T$.
Figure 3: (a,b) Lowest non-negative single-particle Lyapunov spectrum for the measurement-only circuit with (a) PBC and (b) APBC. They are gapped except for $\mu_{e}\simeq 0.5$, where $z_{1}$ behaves as $\sim 1/L$ for PBC and the exact gap closing is anticipated for APBC. (c) Topological invariant after a sufficiently long time, which sharply characterizes topological and trivial phases.
In Fig. 3, we show the lowest single-particle Lyapunov spectrum $z_{1}$ for the measurement-only circuit under the (a) PBC and (b) APBC. It takes a finite value for both cases, implying a finite bulk Lyapunov gap, within each area-law phase. As shown in Fig. 3(c), the topological invariant $\chi_{T}(\bm{s})$ clearly separates the topological and trivial area-law phases. Note that these results are qualitatively independent of the measurement outcomes $\bm{s}$. In fact, $P^{\mathrm{PBC}}_{T}(\bm{s})=1$ is satisfied in the whole parameter region, since the gap does not close at the transition $\mu_{e}=0.5$ for PBC due to finite-size splitting $z_{1}\sim 1/L$. On the other hand, $P^{\mathrm{APBC}}_{T}(\bm{s})$ abruptly changes from $-1$ to $+1$ across the transition point for APBC, implying an exact gap closing near $\mu_{e}=0.5$. This results in the observed transition in $\chi_{T}(\bm{s})$. The behaviors of the bulk Lyapunov gap resemble those of the bulk energy gap $\Delta E$ of the static translation-invariant Kitaev chain [64]; in that case, $\Delta E=0$ for PBC and $\Delta E\sim 1/L$ for APBC at the transition, across which the Pfaffian invariant changes its sign.
The edge-localized zero mode under the OBC in Fig. 2 and the topological invariant in Fig. 3 indicate that the bulk-edge correspondence applies even to nonequilibrium phases under temporally random measurements, if the bulk gap is open. Such an exploration based on topology in random dynamics is accomplished by the Lyapunov analysis and the appropriate definition of the topological invariant via effective Hamiltonians. Note that a naive quantity $P_{T}^{\textrm{PBC}}(\bm{s})P^{\textrm{APBC}}_{T}(\bm{s}^{\prime})$ with $\bm{s}\neq\bm{s}^{\prime}$ does not provide a good topological invariant in general. Rather, focusing on the trajectory under the APBC dependent on the trajectory under PBC makes their fermion parities meaningful to define the invariant under random measurements.
@@
Figure 4: (a) Non-negative single-particle Lyapunov spectrum for the monitored circuit with $J=0.5$ under the OBC for $L=64$. (b,c) Lowest non-negative spectrum $z_{1}$ under the (b) PBC and (c) APBC. (d) Topological invariant at $t=L$ averaged over $1000$ different trajectories. The Lyapunov spectrum and topological invariant distinguish the topological area-law, critical gapless, and trivial area-law phases at $\mu_{e}\simeq 0.4$ and $0.6$, as indicate by the dashed vertical lines.
@@
To discuss the topology of the three phases, we first investigate the Lyapunov spectrum in the circuit with OBC, finding that the critical phase corresponds to a gapless phase without Majorana zero modes. Figure 4(a) shows the $\mu_{e}$ depedence of the non-negative single-particle Lyapunov spectrum. As observed in the measurement-only case, the lowest mode $z_{1}$ become gapless (gapped) in the topological (trivial) phase. In the critical phase for $0.4\lesssim\mu_{e}\lesssim 0.6$, not only the lowest mode $z_{1}$ but also higher modes $z_{k(\neq 1)}$ appear to become gapless, implying the closing of the bulk Lyapunov gap. Interestingly, the $L$-dependence of $z_{1}$ seems different between the topological area-law phase and the critical phase; $z_{1}$ exponentially decays in $L$ in the topological area-law phase ($\mu_{e}\lesssim 0.4$), while it appears to decay in $1/L$ for our available system sizes in the critical phase ($0.4\lesssim\mu_{e}\lesssim 0.6$[53].
@@
Remarkably, however, we find that the topological invariant dynamically characterizes the different phases, including the critical gapless phase. Specifically, we consider the sample-averaged invariant, $\chi_{T}$, for the timescale $T=\mathcal{O}(L)$. Figure 4(d) shows $\chi_{T=L}$ for different system sizes, from which we find three crossings with $\mu_{e}\simeq 0.4,0.5,$ and $0.6$. The locations of the first and last crossings coincide well to the transitions between the area-law gapped phases and critical gapless phase. Notably, this characterization is attributed to the divergent relaxation time of the invariant in the gapless phase. If we denote the timescale as $\tau_{\mathrm{relax}}$, it is at least longer than $\sim z_{1}^{-1}$, which increases faster than $L$ inside the gapless phase. Therefore, $\chi_{T=L}$ will not converge and almost stay zero for the gapless phase. This is contrasted to the gapped phases with $z_{1}=\mathcal{O}(L^{0})$, where $\chi_{T}$ rapidly converge to ±$\pm 1$ for $T=\mathcal{O}(L)$. Note that the crossing point at $\mu_{e}\simeq 0.5$ is different from the boundaries of the three phases. The above argument indicates that this crossing is a finite-size artifact, and $\chi_{T=L}$ will flatly become zero in the entire critical phase for $L\rightarrow\infty$.

### v2

@@ Abstract
We consider a $(1+1)$-dimensional quantum circuit acting on $2L$ Majorana fermions [see Fig. 1(a)], consisting of repeated applications of three elementary steps: (i) unitary time evolution $\hat{\mathcal{U}}$, (ii) measurements of all Majorana pairs on odd bonds, and (iii) measurements of all Majorana pairs on even bonds. These operations update |$\ket{\psi_{t}}$ to |$\ket{\psi_{t+1}}$ in a single time step. To be more precise, the unitary evolution is given by the Kitaev chain Hamiltonian, $\hat{\mathcal{U}}=e^{-i\hat{\mathcal{H}}_{\mathrm{Kitaev}}}$, with
@@
which are weak measurements of the strength $\theta_{e/o}=\tanh^{-1}(\mu_{e/o})$. We set $\mu_{e}=1-\mu_{o}$ hereafter. Given a (unnormalized) state |$\ket{\psi}$, each outcome $s$ is obtained with the Born probability, $p^{e/o}_{j}(s)=\bra{\psi}\hat{\mathcal{K}}_{j}^{e/o}(s)^{\dagger}\hat{\mathcal{K}}_{j}^{e/o}(s)\ket{\psi}/{\langle\psi|\psi\rangle}.$ In the following, we use three boundary conditions: open (OBC), periodic (PBC), and antiperiodic (APBC). For the unitary evolution $\hat{\mathcal{U}}$, the boundary condition is simply specified by $J^{\prime}$: $J^{\prime}=0,J,$ and $-J$ for the OBC, PBC, and APBC, respectively. For the measurements, the OBC means that we discard $\hat{\mathcal{K}}^{e}_{L}(s)$ from the circuit, while we need a careful definition of the APBC, as explained later.
@@
Since both unitary evolution and measurements are bilinear in Majorana fermions, this circuit maps a fermionic Gaussian state to another Gaussian state [46, 47, 48, 49, 50, 51, 52]. A fermionic Gaussian state is completely characterized by the Majorana covariance matrix $\Gamma_{t}$ whose element reads $(\Gamma_{t})_{\ell\ell^{\prime}}=(i/2)\bra{\psi_{t}}[\gamma_{\ell},\gamma_{\ell^{\prime}}]\ket{\psi_{t}}/\langle\psi_{t}|\psi_{t}\rangle$. Time evolution of the (unnormalized) Gaussian state is encoded in the transformation of the Majorana fermions, $\vec{\gamma}\to\hat{\mathcal{K}}_{T}^{\dagger}(\bm{s})\vec{\gamma}(\hat{\mathcal{K}}_{T}^{\dagger}(\bm{s}))^{-1}=K_{T}(\bm{s})\vec{\gamma}$, where
$K_{T}(\bm{s})=\prod_{t=1}^{T}\left(e^{-i\Theta^{e}(\bm{s}_{t})}e^{-i\Theta^{o}(\bm{s}_{t})}e^{H_{\mathrm{Kitaev}}}\right).$ (4)
Here, $H_{\mathrm{Kitaev}}$ and $\Theta^{e/o}(\bm{s}_{t})$ are real antisymmetric matrices defined through ^$\hat{\mathcal{H}}_{\mathrm{Kitaev}}=i\vec{\gamma}^{\mathsf{T}}H_{\mathrm{Kitaev}}\vec{\gamma}/4$ and $\prod_{j}\hat{\mathcal{K}}^{e/o}_{j}(s_{2j/2j-1,t})\propto\exp({i\vec{\gamma}^{\mathsf{T}}\Theta^{e/o}(\bm{s}_{t})}\vec{\gamma}/4)$ with $\vec{\gamma}=(\gamma_{1},\cdots,\gamma_{2L})^{\mathsf{T}}$ (see Supplemental Material [53]).
@@
$H_{\textrm{eff},T}(\bm{s})=-\frac{i}{2T}\ln[K_{T}(\bm{s})K_{T}^{\dagger}(\bm{s})],$ (5)
and the corresponding many-body Hamiltonian ^$\hat{\mathcal{H}}_{\textrm{eff},T}(\bm{s})=-i\vec{\gamma}^{\mathsf{T}}H_{\textrm{eff},T}(\bm{s})\vec{\gamma}/4$. Since $H_{\textrm{eff},T}(\bm{s})$ is real antisymmetric, its eigenvalues $iz_{\ell,T}(\bm{s})\in i\mathbb{R}$ come in pairs, $z_{2j-1,T}(\bm{s})=-z_{2j,T}(\bm{s})$. This defines the single-particle energy spectrum $z_{\ell,T}(\bm{s})$, which in the asymptotic limit coincides with the Lyapunov spectrum: $z_{\ell}=\lim_{T\to\infty}z_{\ell,T}(\bm{s})$. We below arrange them as $0\leq z_{1}=-z_{2}\leq\cdots\leq z_{2L-1}=-z_{2L}$. We can also compute the corresponding Lyapunov vectors $\vec{w}_{\ell,T}(\bm{s})$. Written in the matrix form $\widetilde{O}_{T}(\bm{s})=(\vec{w}_{1,T}(\bm{s}),\cdots,\vec{w}_{2L,T}(\bm{s}))$, they give in the asymptotic limit an orthogonal matrix that brings $H_{\textrm{eff},T}(\bm{s})$ into the standard form,
$\lim_{T\to\infty}\widetilde{O}_{T}^{\mathsf{T}}(\bm{s})H_{\textrm{eff},T}(\bm{s})\widetilde{O}_{T}(\bm{s})=\bigoplus_{j=1}^{L}\begin{pmatrix}{0}&z_{2j-1}\\ -z_{2j-1}&{0}\end{pmatrix}.$ (6)
In practice, we construct $\widetilde{O}_{T}(\bm{s})$ through computing Lyapunov vectors corresponding to the non-negative Lyapunov spectrum based on the complex fermion representation, which is equivalent to the Majorana representation [53]. In the following, the initial state of the trajectories is fixed to be a vacuum state |$\ket{\psi_{0}}=\ket{0}$, whereas the Lyapunov analysis itself is performed with random initial vectors $\vec{w}_{\ell,0}$ for a given trajectory. We only look at a single trajectory specified by a typical sequence of $\bm{s}$ and consider a temporal average of quantities in the long-time limit unless otherwise mentioned [53].
Figure 2: (a) Non-negative single-particle Lyapunov spectra for $L=64$ and (b) sum of squared amplitudes of the spinor wavefunction corresponding to the lowest energy $z_{1}$ at the edges for the measurement-only circuit with the OBC. We find that the topologically non-trivial phase hosts a spatially localized Majorana edge mode with the zero Lyapunov exponent. The inset in (b) shows the spatial distribution of squared amplitudes of the spinor wavefunction at $t=2L$ averaged over $1000$ different trajectories for $L=16$ and $\mu_{e}=0.1$. (c,d) Lowest non-negative single-particle Lyapunov spectrum for the measurement-only circuit with (c) PBC and (d) APBC. They are gapped except for $\mu_{e}\simeq 0.5$, where $z_{1}$ behaves as $\sim 1/L$ for PBC and the exact gap closing is anticipated for APBC. (e) Topological invariant after a sufficiently long time, which sharply characterizes topological and trivial phases.
@@
Remarkably, we reveal that the topological area-law phase is characterized by Majorana zero modes for the Lyapunov spectrum. Figure 2(a) shows the non-negative single-particle Lyapunov spectrum $z_{k}$ against $\mu_{e}$ for the measurement-only circuit under the OBC for $L=64$. Importantly, $z_{1}$ takes a finite value for $\mu_{e}>0.5$, while it vanishes for $\mu_{e}\leq 0.5$. Vanishing of $z_{1}$ under the OBC in the topological area-law phase implies the many-body Lyapunov gap closing for ^$\hat{\mathcal{H}}_{\textrm{eff},T}(\bm{s})$, which is reminiscent of Majorana zero modes for the static Kitaev chain [64].
@@
$P^{\mathrm{PBC/APBC}}_{T}(\bm{s})$ $=\det[\widetilde{O}^{\textrm{PBC/APBC}}_{T}(\bm{s})].$ (9)
Since $\chi_{T}(\bm{s})$ approximates and converges to $Q_{T}(\bm{s})$ for $T\to\infty$, we will also call $\chi_{T}(\bm{s})$ as the topological invariant. We note that $\widetilde{O}^{\textrm{PBC/APBC}}_{T}(\bm{s})$ is not an orthogonal matrix for general $T$.
In Fig. 2(c,d), we show the lowest single-particle Lyapunov spectrum $z_{1}$ for the measurement-only circuit under the (c) PBC and (d) APBC. It takes a finite value for both cases, implying a finite bulk Lyapunov gap, within each area-law phase. As shown in Fig. 2(e), the topological invariant $\chi_{T}(\bm{s})$ clearly separates the topological and trivial area-law phases. Note that these results are qualitatively independent of the measurement outcomes $\bm{s}$. In fact, $P^{\mathrm{PBC}}_{T}(\bm{s})=1$ is satisfied in the whole parameter region, since the gap does not close at the transition $\mu_{e}\simeq 0.5$ for PBC due to finite-size splitting $z_{1}\sim 1/L$. On the other hand, $P^{\mathrm{APBC}}_{T}(\bm{s})$ abruptly changes from $-1$ to $+1$ across the transition point for APBC, implying an exact gap closing near $\mu_{e}=0.5$. This results in the observed transition in $\chi_{T}(\bm{s})$. The behaviors of the bulk Lyapunov gap resemble those of the bulk energy gap $\Delta E$ of the static translation-invariant Kitaev chain [64]; in that case, $\Delta E=0$ for PBC and $\Delta E\sim 1/L$ for APBC at the transition, across which the Pfaffian invariant changes its sign.
The edge-localized zero mode under the OBC in Fig. 2(a,b) and the topological invariant in Fig. 2(e) indicate that the bulk-edge correspondence applies even to nonequilibrium phases under temporally random measurements, if the bulk gap is open. Such an exploration based on topology in random dynamics is accomplished by the Lyapunov analysis and the appropriate definition of the topological invariant via effective Hamiltonians. Note that a naive quantity $P_{T}^{\textrm{PBC}}(\bm{s})P^{\textrm{APBC}}_{T}(\bm{s}^{\prime})$ with $\bm{s}\neq\bm{s}^{\prime}$ does not provide a good topological invariant in general. Rather, focusing on the trajectory under the APBC dependent on the trajectory under PBC makes their fermion parities meaningful to define the invariant under random measurements.
@@
Figure 3: (a) Non-negative single-particle Lyapunov spectrum for the monitored circuit with $J=0.5$ under the OBC for $L=64$. (b,c) Lowest non-negative spectrum $z_{1}$ under the (b) PBC and (c) APBC. (d) Topological invariant at $t=L$ averaged over $1000$ different trajectories. The Lyapunov spectrum and topological invariant distinguish the topological area-law, critical gapless, and trivial area-law phases at $\mu_{e}\simeq 0.4$ and $0.6$, as indicated by the dashed vertical lines.
@@
To discuss the topology of the three phases, we first investigate the Lyapunov spectrum in the circuit with OBC, finding that the critical phase corresponds to a gapless phase without Majorana zero modes. Figure 3(a) shows the $\mu_{e}$ depedence of the non-negative single-particle Lyapunov spectrum. As observed in the measurement-only case, the lowest mode $z_{1}$ become gapless (gapped) in the topological (trivial) phase. In the critical phase for $0.4\lesssim\mu_{e}\lesssim 0.6$, not only the lowest mode $z_{1}$ but also higher modes $z_{k(\neq 1)}$ appear to become gapless, implying the closing of the bulk Lyapunov gap. Interestingly, the $L$-dependence of $z_{1}$ seems different between the topological area-law phase and the critical phase; $z_{1}$ decays faster than power law in the topological area-law phase ($\mu_{e}\lesssim 0.4$), while it appears to decay in $1/L$ for our available system sizes in the critical phase ($0.4\lesssim\mu_{e}\lesssim 0.6$[53].
@@
Remarkably, however, we find that the topological invariant dynamically characterizes the different phases, including the critical gapless phase. Specifically, we consider the sample-averaged invariant, $\chi_{T}$, for the timescale $T=\mathcal{O}(L)$. Figure 3(d) shows $\chi_{T=L}$ for different system sizes, from which we find three crossings with $\mu_{e}\simeq 0.4,0.5,$ and $0.6$. The locations of the first and last crossings coincide well to the transitions between the area-law gapped phases and critical gapless phase. Notably, this characterization is attributed to the divergent relaxation time of the invariant in the gapless phase. If we denote the timescale as $\tau_{\mathrm{relax}}$, it is at least longer than $\sim 1/z_{1}$, which increases faster than $L$ inside the gapless phase. Therefore, $\chi_{T=L}$ will not converge and almost stay zero for the gapless phase. This is contrasted to the gapped phases with $z_{1}=\mathcal{O}(L^{0})$, where $\chi_{T}$ rapidly converge to ±$\pm 1$ for $T=\mathcal{O}(L)$. Note that the crossing point at $\mu_{e}\simeq 0.5$ is different from the boundaries of the three phases. The above argument indicates that this crossing is a finite-size artifact, and $\chi_{T=L}$ will flatly become zero in the entire critical phase for $L\rightarrow\infty$.
@@
Data availability.— Data and simulation codes are available upon reasonable request.